OpenStudy (amistre64):

from a standard deck of playing cards, what is the probability that the second card drawn is a Q of hearts?

OpenStudy (turingtest):

(1/52)(1/51)???? no, can't be...

OpenStudy (anonymous):

If the first card drawn was a Q of hearts, then the probability that the second card drawn is 0

OpenStudy (turingtest):

ha ha, true

OpenStudy (amistre64):

thats correct for one scenario :)

OpenStudy (amistre64):

turning is correct for another scenario; whats the third scenario?

OpenStudy (anonymous):

So we must find the probability that the first card drawn was not a queen of hearts, given that the second card drawn was a queen of hearts!

OpenStudy (turingtest):

if agdgdgwngo is right wouldn't that be (51/52)(1/51) ?

OpenStudy (amistre64):

maybe, but im not quite sure

OpenStudy (anonymous):

Let's run a simulation to find out

OpenStudy (amistre64):

how about this one: what is the probability that the 2nd card is a Qh if the first card is simply drawn and laid to the side without looking at it?

OpenStudy (anonymous):

it is 1/51 * something

OpenStudy (phi):

I would argue that the Q of H is equally likely to be in any of 52 "slots". So the probability of being in a specific slot, e.g. 2nd, is 1/52

OpenStudy (turingtest):

1/51 eh? or is it still 1/52 (which my earlier answer suggests)

OpenStudy (amistre64):

yes, 1/52 is the probability of that last one

OpenStudy (turingtest):

(51/52)(1/51) =1/52 I was right!

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

TuringTest passes the test

OpenStudy (turingtest):

Sweet

OpenStudy (turingtest):

got the idea from you though agdgdgdgwngo

OpenStudy (anonymous):

what idea?

OpenStudy (turingtest):

you said "So we must find the probability that the first card drawn was not a queen of hearts, given that the second card drawn was a queen of hearts!" which made me think the odds of the first card not being QH=51/52, odds of second card being QH=1/51 therefor (51/52)(1/51) like you said